Jeanette is playing in a 9-ball pool tournament. S
Answer #1
45 degree

The original momentum of the cue ball has to be conserved. If you define the original direction of the cue ball as moving along the y-axis, then at the start there is zero momentum along the x-axis and all of it is along the y-axis. After the collision the vector sum of the momentums in the x-direction must still equal zero, meaning the x-component of the velocity of both balls must be equal in magnitude and opposite. Their y-components of velocity must add up to v_i. In addition you know that the angles Theta are equal, so if the x components are equal, Thetas are equal, then the y-components of the velocity must be also equal. You can prove this using v_f Sin(Theta) + v_9 Sin(Theta) = v_i and v_f Cos(Theta) = v_9 Cos(Theta) and conservation of energy (v_f)^2 + (v_9)^2 = (v_i)^2.
Answer #2
solution:

m is the mass of each ball
m*vi=m*(vf*cos?+v9cos?)
and
0=m*(vf*sin?-v9sin?)

and
.5*m*vi^2=.5*m*vf^2+.5*m*v9^2

solve for ?

vi=vf*cos?+v9cos?
0=vf*sin?-v9sin?
vf=v9

vi^2=2*vf^2
vi=2*vf*cos?

vi^2=4*vf^2*cos^2?

sqrt(.5)=cos?
?=45 degrees